OpenStudy (anonymous):

Find the negative square root of 48

OpenStudy (breanna1999):

Any answer choices?

OpenStudy (anonymous):

no

OpenStudy (breanna1999):

Hmm the one I got is -6.9

OpenStudy (anonymous):

Ok do you know how you would get that without using a calclutor?

OpenStudy (breanna1999):

I do not

OpenStudy (anonymous):

ok thanks

OpenStudy (phi):

you could factor 48 into its prime factors 2*2*2*2 * 3 and look for "pairs" there are two pairs of "2's". "pull each pair out of the radical, and put only one of each out front: \[ \sqrt{2\cdot 2 \cdot 2 \cdot 2 \cdot 3}= 2\cdot 2 \sqrt{3} \] simplify 2*2 to 4 and we want the negative square root, so make it negative \[ -4 \sqrt{3} \] notice what we get if we multiply it by itself \[ -4 \sqrt{3} \cdot -4 \sqrt{3} = -4\cdot -4 \cdot \sqrt{3} \cdot \sqrt{3} \\ =16 \cdot 3 = 48 \]

OpenStudy (anonymous):

ok thanks i think i understand now